# Highly efficient quantum Stirling engine using multilayer Graphene

### v2

@@ IV.3 Stirling Operational Regimes
This does not exclude bilayer graphene from exhibiting the full range of operational regimes. Previous studies have shown that the same Stirling cycle implemented in bilayer graphene can display all operational regimes under different parameter choices and when the chemical potential is not fixed Soufy and Benjamin (2025). Their absence here therefore reflects the specific thermodynamic window considered, rather than an intrinsic limitation of the system.
@@ V Conclusions and discussion
In this study, quantum Stirling cycles in monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene under perpendicular magnetic fields were investigated within a grand-canonical, fully fermionic framework based on the exact Landau spectra. An order-of-magnitude comparison between the particle number $N_{e}$, the chemical potential employed in the simulations, and sheet carrier densities defined by the magnetic length indicates that the most favorable regimes are experimentally accessible.
@@ Acknowledgements.
B.C, F.J.P., M.H.G, N.C, and P.V. acknowledge financial support from ANID Fondecyt grant no. 1240582. B.C, F.J.P., M.H.G, and P.V. acknowledge financial support from ANID Fondecyt grant no. 1250173. B.C acknowledge PUCV and Programa de Incentivo a la Iniciación Científica (PIIC) no. 004 from ”Dirección de Postgrado” of UTFSM. B.C. acknowledges the support of ANID Becas/Doctorado Nacional 21250015. C.L. thanks the support by the Brazilian agencies CNPq and FAPERJ.
@@ A.1 Efficiency and work at fixed $\mu$
Figure 10: Stirling-cycle efficiency $\eta$ of AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=19$, $20$, and $21~\mathrm{meV}$. The cycle parameters are the same as those used in the main text: $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$, corresponding to a Carnot efficiency $\eta_{C}=0.037$ (horizontal dashed line). The vertical red line marks the value of $B_{L}$ at which the fixed-particle-number protocol with $N_{e}=5$ yields $\eta=\eta_{C}$ in the main analysis.
@@
Figure 10 shows the efficiency $\eta$ as a function of the lower magnetic field $B_{L}$ for several representative values of $\mu$, using the same cycle parameters as in the main text, $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$, for which the Carnot efficiency is $\eta_{C}=0.037$. The vertical red line indicates the value of $B_{L}$ at which the fixed-particle-number protocol with $N_{e}=5$ attains $\eta=\eta_{C}$.
Figure 11: Total work per cycle $W$ for AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=15$, $20$, and $38~\mathrm{meV}$. The cycle parameters are $T_{L}=10~\mathrm{K}$, $T_{H}=50~\mathrm{K}$, and $B_{H}=1.5~\mathrm{T}$. The oscillatory behavior originates from successive Landau-level crossings through the fixed chemical potential, which modulate the electronic population and the heat exchanged during the isothermal strokes.
@@
Figure 12: Stirling-cycle efficiency $\eta$ of AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=15$, $20$, and $38~\mathrm{meV}$, using the same cycle parameters as in Fig. 11. The horizontal dashed line indicates the Carnot efficiency $\eta_{C}$. Despite the oscillatory features induced by Landau-level crossings, well-defined high-efficiency operating regimes persist.
@@ A.2 Particle-number variation at fixed $\mu$
To further clarify the relation between the fixed-$N_{e}$ and fixed-$\mu$ formulations, Fig. 13 presents a contour plot of the particle number $\langle N\rangle$ as a function of temperature $T$ and magnetic field $B$ at fixed chemical potential $\mu=19~\mathrm{meV}$. The parameters correspond to the same range used in the main text, and the vertical line marks the value of $B_{L}$ associated with the Carnot point identified in the fixed-$N_{e}$ analysis for $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$.
@@
Figure 13: Contour plot of the particle number $\langle N\rangle$ as a function of temperature $T$ and magnetic field $B$ for AB-stacked bilayer graphene at fixed chemical potential $\mu=19~\mathrm{meV}$. The parameter range corresponds to that used in the main text. The vertical line marks the value of $B_{L}$ associated with the Carnot point identified in the fixed-$N_{e}$ analysis for $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$. In this region, the nearly horizontal contour lines indicate that the particle number varies only weakly with temperature, explaining the consistency between the fixed-$N_{e}$ and fixed-$\mu$ formulations in the operating regime of interest.

### v3

@@ IV.3 Stirling Operational Regimes
This does not exclude bilayer graphene from exhibiting the full range of operational regimes. Previous studies have shown that the same Stirling cycle implemented in bilayer graphene can display all operational regimes under different parameter choices and when the chemical potential is not fixed [56]. Their absence here therefore reflects the specific thermodynamic window considered, rather than an intrinsic limitation of the system.
@@ V Conclusions and discussion
This study investigates quantum Stirling cycles in monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene subjected to perpendicular magnetic fields. Using a grand-canonical, fully fermionic framework based on exact Landau spectra, we demonstrate that the cycles can operate within experimentally accessible regimes. This is supported by an order-of-magnitude comparison between simulated particle numbers, chemical potentials, and sheet carrier densities defined by the magnetic length.
@@ Acknowledgements.
B.C, F.J.P., M.H.G, N.C, and P.V. acknowledge financial support from ANID Fondecyt grant no. 1240582. B.C, F.J.P., M.H.G, and P.V. acknowledge financial support from ANID Fondecyt grant no. 1250173. B.C acknowledge PUCV and Programa de Incentivo a la Iniciación Científica (PIIC) no. 004 from ”Dirección de Postgrado” of UTFSM. B.C. acknowledges the support of ANID Becas/Doctorado Nacional 21250015. C.L. thanks the support by the Brazilian agencies CNPq and FAPERJ. M.H.G acknowledge PUCV and ”Dirección de Postgrado” of UTFSM.
@@ A.1 Efficiency and work at fixed $\mu$
Figure 10: Stirling-cycle efficiency $\eta$ of AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=19$, $20$, and $21~\mathrm{meV}$. The cycle parameters are the same as those used in the main text: $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$, corresponding to a Carnot efficiency $\eta_{C}=0.037$ (horizontal dashed line). The vertical red line marks the value of $B_{L}$ at which the fixed-particle-number protocol with $N_{e}=5$ yields $\eta=\eta_{C}$ in the main analysis. For values exceeding $\eta_{C}$ or falling below zero, the system transitions to a different operational regime.
@@
Fig. 10 shows the efficiency $\eta$ as a function of the lower magnetic field $B_{L}$ for several representative values of $\mu$, using the same cycle parameters as in the main text, $T_{L}=45.9~\mathrm{K}$, $T_{H}=47.7~\mathrm{K}$, and $B_{H}=3.18~\mathrm{T}$, for which the Carnot efficiency is $\eta_{C}=0.037$. The vertical red line indicates the value of $B_{L}$ at which the fixed-particle-number protocol with $N_{e}=5$ attains $\eta=\eta_{C}$. For values below zero and above $\eta_{C}$, the definition of efficiency is no longer applicable because the system transitions to a different operational regime. To avoid possible confusion, these regions are therefore not plotted.
Figure 11: Total work per cycle $W$ for AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=15$, $20$, and $38~\mathrm{meV}$. The cycle parameters are $T_{L}=10~\mathrm{K}$, $T_{H}=50~\mathrm{K}$, and $B_{H}=1.5~\mathrm{T}$. The oscillatory behavior originates from successive crossings of the chemical potential through the Landau levels, which modulate the electronic population and the heat exchanged during the isothermal strokes. For value falling below zero, the system transitions to a different operational regime.
@@
Figure 12: Stirling-cycle efficiency $\eta$ of AB-stacked bilayer graphene as a function of the lower magnetic field $B_{L}$, computed at fixed chemical potential $\mu=15$, $20$, and $38~\mathrm{meV}$, using the same cycle parameters as in Fig. 11. The horizontal dashed line indicates the Carnot efficiency $\eta_{C}$. Despite the oscillatory features induced by the crossings of the chemical potential through the Landau levels, well-defined high-efficiency operating regimes persist. For values falling $\eta_{C}$ below zero, the system transitions to a different operational regime.
@@ A.2 Particle-number variation at fixed $\mu$
To further clarify the relation between the fixed-$N_{e}$ and fixed-$\mu$ formulations, Fig. 13 presents a contour plot of the particle number $N_{e}$ as a function of temperature $T$ and magnetic field $B$ at fixed chemical potential $\mu=19~\mathrm{meV}$. The parameters correspond to the same range used in the main text.
@@
Figure 13: Contour plot of the particle number $N_{e}$ as a function of temperature $T$ and magnetic field $B$ for AB-stacked bilayer graphene at fixed chemical potential $\mu=19~\mathrm{meV}$. The parameter range corresponds to that used in the main text. In this region, the nearly horizontal contour lines indicate that the particle number varies only weakly with temperature, explaining the consistency between the fixed-$N_{e}$ and fixed-$\mu$ formulations in the operating regime of interest.